# AI Hospital: Interactive Evaluation and Collaboration of LLMs as Intern Doctors for Clinical Diagnosis

### v2

@@ Abstract
The incorporation of Large Language Models (LLMs) in healthcare marks a significant advancement. However, the application has predominantly been limited to discriminative and question-answering tasks, which does not fully leverage their interactive potential. To address this limitation, our paper presents AI Hospital, a framework designed to build a real-time interactive diagnosis environment. To simulate the procedure, we collect high-quality medical records to create patient, examiner, and medical director agents. AI Hospital is then utilized for the interactive evaluation and collaboration of LLMs. Initially, we create a Multi-View Medical Evaluation (MVME) benchmark where various LLMs serve as intern doctors for interactive diagnosis. Subsequently, to improve diagnostic accuracy, we introduce a collaborative mechanism that involves iterative discussions and a dispute resolution process under the supervision of the medical director. In our experiments, we validate the reliability of AI Hospital. The results not only explore the feasibility of apply LLMs in clinical consultation but also confirm the effectiveness of the dispute resolution focused collaboration method. [^1] https://github.com/LibertFan/AI_Hospital
@@ 1 Introduction
The realm of healthcare has witnessed a remarkable evolution with the integration of artificial intelligence Bajwa et al. (2021), particularly through the development of Large Language Models (LLMs) OpenAI (2023); Touvron et al. (2023); Bai et al. (2023). These advancements open novel pathways for medicine Singhal et al. (2023); Chen et al. (2023e); Bao et al. (2023); Zhang et al. (2023), with LLMs demonstrating remarkable performance across various medical question-answering benchmarks Jin et al. (2021, 2019); Pal et al. (2022). Nonetheless, the complexity of real-world medical consultations, which typically involve a series of steps such as assessing patient conditions, performing physical and auxiliary examinations, and offering medication and health guidance, presents a challenge. These steps require intricate, multi-turn interactions with patients, raising questions about the efficacy of LLMs in managing real-world medical consultations.
In this work, we introduce AI Hospital, a platform based on multi-agent interaction Park et al. (2023); Li et al. (2023) to build a real-time interactive consultation scenario, and explore the feasibility of using LLMs as doctors in clinical diagnosis. The platform incorporates various roles including intern doctors, patients, examiners, and a medical director. In the framework, intern doctors diagnose diseases through active questioning, patients need to describe their symptoms and respond to inquiries from doctors, examiners execute medical tests following the doctors’ instructions, while the medical director is tasked with assessing the performance of the intern doctors in diagnosis. Significantly, the primary objective of AI Hospital is twofold: firstly, facilitating interactive consultation among intern doctors, patients and examiners to evaluate the diagnostic capability of LLMs, and secondly, enabling collaborative efforts among doctors, thereby improving the diagnostic accuracy.
Figure 1: The demonstration of AI Hospital.
Establishing AI Hospital necessitates overcoming the challenge of replicating the behavior of personnel in medical settings, such as patients and examiners. (1) To address the issue of medical roles lacking medical knowledge, we collect high-quality Chinese medical records, utilizing professional medical data as inputs for large language models to effectively simulate the agents. (2) Given the critical role of professional medical examinations in clinical diagnosis, with CDC finding that 70% medical decisions rely on them di Ruffano et al. (2012); CDC (2018), we establish an examiner role specifically for recording and overseeing patients’ physical and auxiliary examination results. This configuration also restricts patient agents from accessing detailed medical examination data, ensuring the alignment in behavior with actual patients. (3) To mitigate patients’ lack of self-awareness in conversation scenarios Wang et al. (2023c), we craft background narratives, personas and descriptions of conditions with personal characteristics to guide the conversational style. In summary, we intricately craft the prompts for these medical roles, allowing for natural interaction with intern doctors grounding on the setting of role.
Based on the AI Hospital, we investigate the feasibility of utilizing LLMs as intern doctors for clinical diagnosis by establishing a Multi-View Medical Evaluation (MVME) benchmark. The benchmark assesses the real-time interactive consultation capabilities of LLMs across three critical dimensions. Firstly, it examines their ability to identify patient symptoms, highlighting the importance of actively seeking relevant information. Secondly, we assess the comprehensiveness of their medical examinations, specifically their adeptness in selecting and administering a suitable range of these tests. Thirdly, we measure the accuracy and professionalism of their diagnosis, checking whether they meet standards of medical practice. In practice, we not only leverage GPT-4 as a flexible evaluation engine but also employ a link-based method to incorporate medical standard knowledge.
@@
To enhance LLMs’ diagnostic accuracy, drawing on previous research Croft et al. (2015); Llewelyn et al. (2014); Centor et al. (2019) that highlights the importance of teamwork in clinical diagnosis, we delve into the collaborative mechanism O’leary et al. (2010); Lamb et al. (2011), facilitating iterative discussions among intern doctors. Different from previous models that lack concise organization and clear direction for participants Talebirad and Nadiri (2023); Liang et al. (2023), our framework introduces a dispute resolution strategy that effectively engages the medical director to guide discussions, clarify issues, and steer the interns towards a more structured and efficient collaboration, thereby fostering more focused discussions and accelerating the achievement of consensus.
In our experiment, we initially investigate the behavior of various personnel within the AI Hospital to validate the reliability. Subsequently, we test a range of LLMs in the AI Hospital, uncovering that deficiencies in interaction significantly impair diagnostic accuracy. Furthermore, we assess the effectiveness of our dispute resolution-focused cooperative mechanism, demonstrating its capacity to accelerate and enhance diagnostic accuracy. Finally, we dive into GPT-4, analyzing its limited efficacy in identifying necessary medical examinations and managing complications.
In general, our contributions are of three-folds:
  • We develop AI Hospital underpinned by an LLM-driven multi-agent framework.
  • We build a benchmark MVME for interactive evaluation within the medical domain to assess LLMs in interactive consultations.
  • We design a collaborative mechanism which focuses on dispute resolution. It demonstrates diagnostic capabilities surpassing GPT-4.
@@ 2 Setup of AI Hospital
The framework of our AI Hospital is illustrated in Figure 1. It encompasses three principal procedures: (1) Intern doctors proactively question patients about their physical condition and patients obtain diagnostic reports from intern doctors; (2) Patients undertake medical examinations following the directions of intern doctors; (3) Medical director assess the performance of intern doctors. In this section, §2.1 details how we utilize the collected medical records to equip medical roles with necessary medical expertise. Subsequently, §2.2 to §2.5 describe the methodologies employed to establish the four agents, including the intern doctor.
2.1 Agents Setup on Medical Records
We collect a substantial volume of Chinese medical record data, comprising various sections. Given the requirement of specific medical knowledge in different roles, we categorize the information of medical record into three types.
Basic Information. It is consisting of chief complaints, current and past medical histories, and personal and familial histories.
Professional Medical Examination. It encompasses physical and auxiliary examinations.
Diagnosis and Treatment. It has diagnostic results, diagnostic rationales, and treatment course.
These categories are correspondingly allocated to the patient, examiner, and medical director, then employed in prompts of LLMs to facilitate the medical knowledge of distinct roles.
2.2 Intern Doctor
In AI Hospital, intern doctors are responsible for clinical diagnosis. This requires them to actively engage with patients, delving into detailed inquiries about their physical conditions. Through these interactions, intern doctors are expected to apply their medical knowledge and skills to identify potential illnesses afflicting the patients. Subsequently, the doctors conduct medical examinations through examiners to carry out a thorough investigation. Once a diagnosis is identified, the intern doctors then move to the formulation of treatment plans. They are expected to consider various treatment options, weigh their effectiveness, and select the most appropriate course of action. The prompt of intern doctor is shown in Table 12.
Figure 2: The workflow of examiner.
2.3 Examiner
In an effort to replicate the functions of a hospital’s examination department, we introduce the examiner agent for the administration of medical examination. It is designed to manage “Professional Medical Examination” as detailed in § 2.1. When patient agents ask for specific tests under physician guidance, the examiner agent should provide the relevant examination results.
To enhance the rationality of the inspection process and ensure that results are provided only upon receipt of correct examination items, we divide the examination procedure into two sequential stages as shown in Figure 2. In the first stage, examiners are tasked with accurately interpreting patients’ examination requests and discerning the required tests as the prompt in Table 10. This involves extracting essential examination items and dismissing requests with vague objectives, such as “inquiries about general medical examination conditions”. In the second phase, examiners follow the prompt in Table 11 to convey the examination outcomes to the doctors, contingent upon the examination items identified in the initial phase.
2.4 Patient
Functionality. We conduct a systematic examination of the essential capabilities required from patients during interactive diagnosis. These capabilities include: (1) the articulate expression of complaints; (2) truthful responses to doctors’ inquiries; (3) active participation in examinations; and (4) the voicing of questions and concerns. We tailor prompts accordingly to meet the requirements. The prompt is shown in Table 9.
Personality. To facilitate improved behavior control, we propose enhancing self-awareness in patient role-playing through the development of a patient’s personality. First, we utilize GPT-4 to reason and imagine the detailed aspects of the patient’s lifestyle, pertinent events, personality traits, and speaking manner that correlate with their physical condition. There is an example: “You are a farmer who left school after primary education. Your life is frugal, but your mindset is optimistic. You lack knowledge of medical terminology.” Second, in line with the patient’s profile, we adopt GPT-4 to generate a statement reflecting their condition. For example, “Doctor, I’ve been having this pain on the right side of my stomach, and I don’t want to eat at all. Also, I’ve been peeing blood these past couple of days.” This statement serves as the first segment of the interactive consultation for one-shot demonstration. By crafting comprehensive patient profiles as prompts and employing one-shot learning, the patient agent can replicate the behaviors of a real patient in dialogues.
2.5 Medical Director
The primary responsibility of the medical director is to assess the performance of intern doctors in interactive diagnosis. At the end of a consultation, the medical director asks the intern doctors to provide a comprehensive summary that includes the patient’s symptom, medical examinations, diagnostic results, diagnostic rationales, and the treatment plan. This procedure is designed to consolidate information regarding the patient’s state, elucidate the interns’ comprehension of the diagnosed condition, the basis of their judgments, and the treatment strategy. Upon receiving responses about these five areas, the medical director evaluates each section by comparing it with the medical record, which serves as the gold standard. For more detailed information of evaluation, please refer to §3.
Figure 3: Collaboration of intern doctors for clinical diagnosis.
@@ 3 MVME: Evaluation of LLMs as Intern Doctors for Clinical Diagnosis
Based on AI Hospital, we assess the feasibility of employing LLMs for clinical diagnosis.
@@ 3.1 Multi-View Evaluation Criteria
We develop evaluation criteria from three aspects to assess the procedure of clinical diagnosis.
Mastery of Patient Symptoms. This assesses how precisely the intern doctors can gather and understand information about patient symptoms. It focuses on their ability to comprehensively explore the medical history, identify critical symptoms based on patient descriptions, and perform necessary analyses and documentation.
Comprehensiveness of Medical Examination. This evaluates the range and appropriateness of the medical tests employed by the interns. It considers their choice and skill in interpreting the findings.
Quality of Diagnostic Report. This measures the accuracy and professionalism of the diagnostic report, including the diagnostic results, diagnostic rationales, and the proposed treatment plan. It examines the report’s consistency with medical logic and standards.
Centering on three evaluation criteria, our assessment covers five specific aspects: (1) symptom, (2) medical examination, (3) diagnostic results, (4) diagnostic rationales, and (5) treatment plan.
3.2 Evaluation Methods
LLM based Evaluation with Reference
To assist the medical director in evaluation of the five specific portions, we extract the “Diagnosis and Treatment” with “Chief Complaints” and “Auxiliary Examinations” from the medical records which are described in §2.1. The information is incorporated into the prompts for the LLMs (GPT-4) as Table 13&14, providing external expertise for the medical director to assess LLMs. In each specific evaluation dimension, medical director rate the interns’ performance on a scale of 1 to 4, ranging from poorest to excellent.
Link based Evaluation of Diagnostic Results
To enhance the scientific rigor and validity of our evaluations, particularly in diagnostic results with respect to identify diseases, we employ the International Classification of Diseases (ICD-10) Trott (1977) as the authoritative source, and link standardized disease terminologies with natural language based diagnostic results. Initially, we extract disease entities from the diagnostic results and the medical records’ references. Subsequently, we implement a fuzzy matching process with a predefined threshold of 0.5 to link these disease entities with ICD-10 terminology, building a normalized disease sets. Ultimately, we assess diagnostic agreement between the diagnoses from intern doctors and the golden reference, then report the number of extracted disease entities (#), set-level precision (P), recall (R), and F1-score (F) metrics.

### v3

# AI Hospital: Benchmarking Large Language Models in a Multi-agent Medical Interaction Simulator

@@ Abstract
Artificial intelligence has significantly advanced healthcare, particularly through large language models (LLMs) that excel in medical question answering benchmarks. However, their real-world clinical application remains limited due to the complexities of doctor-patient interactions. To address this, we introduce AI Hospital, a multi-agent framework simulating dynamic medical interactions between Doctor as player and NPCs including Patient, Examiner, Chief Physician. This setup allows for realistic assessments of LLMs in clinical scenarios. We develop the Multi-View Medical Evaluation (MVME) benchmark, utilizing high-quality Chinese medical records and NPCs to evaluate LLMs’ performance in symptom collection, examination recommendations, and diagnoses. Additionally, a dispute resolution collaborative mechanism is proposed to enhance diagnostic accuracy through iterative discussions. Despite improvements, current LLMs exhibit significant performance gaps in multi-turn interactions compared to one-step approaches. Our findings highlight the need for further research to bridge these gaps and improve LLMs’ clinical diagnostic capabilities. Our data, code, and experimental results are all open-sourced at https://github.com/LibertFan/AI_Hospital.
@@ 1 Introduction
Healthcare has seen significant progress with artificial intelligence in recent years [1], especially through the development of large language models (LLMs) [2, 3, 4]. These models demonstrate impressive performance on static medical question answering benchmarks like MedQA [5], PubMedQA [6], and MedMCQA [7], even rivaling human experts. However, a significant gap remains between LLMs performance on these benchmarks and their real-world application in clinical diagnosis.
In practice, patients frequently lack sufficient medical knowledge and may possess ambiguous understandings of various medical concepts, making it difficult for them to accurately and comprehensively communicate their physical condition to healthcare providers in a single interaction [8]. As a result, patients often exhibit a strong dependence on doctors for guidance and clarification. Doctors assume a leading role in the complex, multi-turn interactions with patients to gather the information required for accurate diagnosis and treatment [9, 10]. Despite the crucial importance of this dynamic, patient-centered diagnostic process, there is a scarcity of research assessing the ability of large language models (LLMs) to simulate this interaction.
Figure 1: The demonstration of AI Hospital framework.
To address above challenges, we introduce AI Hospital, a LLM-powered multi-agent framework that simulates real-world dynamic medical interactions. AI Hospital consists of multiple non-player characters (NPCs), including Patient, Examiner, and Chief Physician, as well as the player character, represented by the Doctor. It recreates the scenario of a Patient visit, requiring Doctor to engage in multi-turn conversations with Patient, ask relevant and probing questions, recommend appropriate medical examinations, and make diagnoses after collecting enough information. We set up Examiner in the NPCs, who are specifically tasked with interacting with Patient and providing pertinent medical examination results, ensuring that Doctor have access to the necessary objective information of the Patient to make accurate diagnoses. Additionally, Chief Physician is responsible for evaluating the performance of Doctor after the entire session. The multi-agent nature of the AI Hospital framework allows for realistic simulations of complex medical scenarios, enabling a comprehensive assessment of an LLMs’ ability to navigate various clinical situations.
Based on the AI Hospital framework, we investigate the feasibility of utilizing LLMs as Doctors for clinical diagnosis by establishing the Multi-View Medical Evaluation (MVME) benchmark. This benchmark incorporates a collection of high-quality Chinese medical records meticulously screened by experienced medical professionals. These real-world cases provide detailed structured medical profiles, including patient subjective conditions, objective medical test results, diagnoses and treatments. Leveraging GPT-3.5 and GPT-4 to simulate the AI Hospital’s non-player characters (NPCs), we conduct a thorough evaluation of the performance of LLMs-driven Doctors within this dynamic and realistic medical interaction environment. The MVME benchmark evaluates the performance of Doctor along three key dimensions by Chief Physician: the ability to collect symptoms, recommend examinations, and make diagnoses. As a supplement, we also develop a link-based approach that integrates medical standard knowledge to generate entity-level evaluation metrics.
@@
To enhance LLMs’ diagnostic accuracy, drawing on previous research [11, 12] that highlights the importance of teamwork in clinical diagnosis, we delve into the collaborative mechanism [13, 14], facilitating iterative discussions among doctors. We utilize multiple Doctors to independently engage with the same medical record, allowing for diverse conversation trajectories and diagnostic outcomes. A dispute resolution strategy is proposed that effectively engages a Centre Agent to guide discussions, clarify issues, and steer the Doctors towards a more structured and efficient collaboration, thereby fostering more focused discussions and accelerating the achievement of consensus.
To assess the performance of LLMs in clinical diagnosis, we conduct extensive experiments within the AI Hospital framework. We first validate the reliability of the various roles in AI Hospital and then evaluate a range of LLMs in the interactive diagnostic process. Our results reveal a substantial performance gap between interactive LLMs and the one-step GPT-4, which serves as an upper bound by directly utilizing all patient information in a single interaction. On key metrics such as the accuracy of diagnostic results, diagnostic rationales, and treatment plans, the performance of interactive LLMs is less than 50% of that achieved by the one-step GPT-4. Despite dedicated prompting engineering, LLMs struggle to make reasonable decisions in multi-turn interactions, leading to suboptimal diagnostic accuracy. The dispute resolution collaborative mechanism enhances performance to a certain degree but also falls short of the upper bound. It’s suggested that existing LLMs may not have fully assimilated effective multi-turn diagnostic strategies. Our quantitative results highlights the challenges existing LLMs face in posing pertinent questions, eliciting crucial symptoms, and recommending appropriate medical examinations. These findings underscore the difficulties encountered by current LLMs in replicating the complex clinical reasoning processes employed by professional doctors and emphasize the need for further research to bridge the gap between LLMs and human physicians in clinical diagnosis.
In summary, the main contributions of this paper can be summarized as follows: 1) We introduce AI Hospital, a novel LLM-powered multi-agent framework to simulate medical interactions, enabling comprehensive evaluation of LLMs’ ability to navigate complex clinical scenarios; 2) We establish the Multi-View Medical Evaluation (MVME) benchmark, which leverages high-quality medical records to evaluate the performance of LLMs-driven doctors in collecting symptoms, recommending examinations, and making diagnoses; 3) We propose a dispute resolution collaborative mechanism that facilitates iterative discussions among doctors to enhance diagnostic accuracy. The potential of our AI Hospital framework is comprehensively discussed in Appendix H.
@@ 2 Setup of AI Hospital
As depicted in Figure 1, the AI Hospital framework comprises three NPC characters — the Patient, Examiner, and Chief Physician — and one player character, the Doctor. Each character assumes specific roles and responsibilities within the framework. The AI Hospital operates in two phases. In the diagnostic phase, Patient, Examiner, and Doctor engage in conversations to exchange information necessary for accurate diagnosis. The number of interaction turns in this phase can vary depending on the Doctor’s diagnostic strategy. Subsequently, during the evaluation phase, Chief Physician is responsible for scoring the performance of Doctor in the diagnostic phase. The following sections will elaborate on the settings and construction methods for each agent in AI Hospital.
2.1 Agents Setup with Medical Records
Medical records are a valuable resource for reconstructing the hospital visit experience and simulating real-world medical interactions. By leveraging these records, we can reverse-engineer the diagnostic process and shape the behavior of agents within the AI Hospital framework. We categorize the information in each medical record into three types: 1) Subjective Information  This category includes the patient’s symptoms, etiology, past medical history, habits, etc., which are primarily provided by the patient during their verbal interactions with the doctor; 2) Objective Information  This category encompasses medical test reports such as complete blood counts, urinalysis, and chest X-rays. The presence of these data in medical records indicates that the patient underwent these tests during the diagnostic process at the doctor’s recommendation; 3) Diagnosis and Treatment  This category consists of diagnostic results, diagnostic rationales, and treatment courses, which are the final conclusions made by the doctor during the diagnostic process, based on the combination of subjective and objective information.
These categories of information are assigned to the corresponding agents in the AI Hospital framework. Patient has access to the subjective information, Examiner is aware of the objective information, and Chief Physician possess all information, while Doctor do not have access to any information. AI Hospital framework assigns specific categories of information from medical records to each agent, shaping their scope of information within the diagnostic process.
2.2 Agent Behavior Setting for NPCs
In the AI Hospital framework, we leverage GPT-3.5 to power Patient and Examiner, and GPT-4 to drive Chief Physician, enabling them to embody their roles authentically. Beyond providing NPCs with relevant information in medical records, we also employ meticulous prompt engineering to encourage they exhibit realistic behavior patterns.
Patient  The Patient agent is designed to exhibit a set of realistic behavior patterns to enhance the authenticity of the medical simulation: 1) Cooperation. The agent should actively respond to the doctor’s inquiries and provide truthful answers, even if they may not proactively disclose all relevant information. The agent should actively participate in medical examinations recommended by the doctor; (2) Communication. The agent should use colloquial language and may omit certain important details or have subjective biases in describing their condition due to limited medical knowledge or personal beliefs; 3) Curiosity. The agent should express concerns or questions based on their level of understanding, seeking clear explanations from the doctor to address their doubts about the diagnosis or treatment process; 4) Personalisation. For each medical record, we employ GPT-4 to reason and imagine the patient’s unique background, experiences, emotional responses, and personality traits, thereby enhancing the realism and depth of the simulation. The prompt for Patient agent is shown in Table 9.
Examiner  The Examiner agent’s primary goal is to provide relevant examination results when the Patient agent requests a query for a specific medical test. To maintain the authenticity of the simulation, the agent follows a realistic workflow. Upon receiving an examination query, the agent first identifies the requested medical examination and rejects any ambiguous or unclear requests. If the corresponding medical examination results are available, the Examiner agent returns the relevant findings to the doctor. In cases where no specific results are found, the agent reports no abnormalities. The Prompts are shown in Table 11 and  11.
Chief Physician  The primary responsibility of the Chief Physician agent is to evaluate the performance of the Doctor agent in interactive diagnosis. After the diagnostic phase, Chief Physician first requires the Doctor to provide a comprehensive summary report for the patient, and then evaluates the summary report by comparing it with raw medical record, which serves as the gold standard. The prompts for Chief Physician agent and more detailed description of the evaluation process can be found in Table 13 and § 3 respectively.
2.3 Agent Behavior Setting for Player
The player agent, i.e., the Doctor, can be powered by various LLMs that are being evaluated. However, in order to be able to engage in conversations based on predefined settings, LLMs are required to be well instruction-followed, otherwise LLMs will struggle to interact in AI Hospital.
Doctor  The Doctor agent is designed to emulate the essential qualities and duties of a skilled and empathetic physician in real-world practice. The agent is encouraged to actively gather information, focusing on obtaining the patient’s physical conditions like symptoms and medical history. When the agent determines that additional objective data is necessary to arrive at a confident diagnosis or to confirm a suspected condition, it suggests relevant examinations and tests. By synthesizing both subjective and objective findings, the agent aims to accurately diagnose the patient’s condition, mirroring the systematic approach used by experienced doctors. The prompt of intern doctor is shown in Table 15.
2.4 Dialogue Flow in AI Hospital
The AI Hospital framework simulates a realistic diagnostic process through a structured dialogue flow involving multiple agents. The conversation is initiated by the Patient agent, who presents a chief complaint generated by GPT-4 based on the patient’s medical record. The Doctor agent then engages in a series of interactions with the Patient and Examiner agents to gather necessary information and make an accurate diagnosis. Throughout the dialogue, each agent’s responses are prefixed with special symbols to explicitly indicate the intended recipient of their message, enabling a seamless multi-party conversation flow. The dialogue continues until the Doctor agent reaches a diagnosis or a predefined maximum number of interaction rounds is reached. For a more detailed description of the dialogue flow, please refer to Appendix A.
@@ 3 MVME: Evaluation of LLMs as Intern Doctors for Clinical Diagnosis
Based on AI Hospital, we assess the feasibility of employing LLMs as Doctor agent for clinical diagnosis by establishing the Multi-View Medical Evaluation (MVME) benchmark.
@@ 3.1 Multi-View Evaluation Criteria
Evaluating the performance of the Doctor agent is a crucial component of the AI Hospital framework. As mentioned in § 2.2, in the evaluation phase, the Doctor is required to provide a comprehensive summary report of the Patient. We require the summary report consists of 5 parts, including the patient’s symptoms, medical examinations, diagnostic results, diagnostic rationales, and treatment plan.
Since the contents of the patient’s medical record are described using natural language, the Chief Physician, as the evaluator, will directly compare each part of the report with the patient’s complete medical record. For each part of the summary report, the GPT-4-driven Chief Physician needs to score from four discrete scores: 1, 2, 3, and 4, representing poorest to excellent performance. The evaluation of the "symptoms" part can reflect the comprehensiveness of the symptoms collected by the Doctor during the interaction process. The evaluation of the "medical examinations results" part can reflect the appropriateness of the medical examinations suggested by the Doctor. The evaluation of the other parts can reflect the Doctor’s diagnostic and treatment capabilities. These metrics can reflect the LLMs’ both dynamic and static medical decision-making abilities, including proactive inquiry, information gathering, clinical knowledge and comprehensive judgment.
In addition to the above model-based evaluation method, we also compute entity-overlap-based automated metrics for the diagnostic results part. We extract all disease entities from the diagnostic results provided by the LLMs and the actual medical records, and link them to their corresponding standardized disease entities. We then calculate the entity overlap to measure the accuracy of the final diagnoses made by the LLMs. We report the average number of extracted disease entities (#), set-level precision (P), recall (R), and F1 score (F) metrics. Currently, we only calculate entity-level metrics for diseases because readily available entity linking methods and standards, such as the International Classification of Diseases (ICD-10) [15], exist for this purpose. For symptoms and examinations, linking them to corresponding entities is more challenging due to the lack of readily available tools. Therefore, we do not incorporate their entity-level metrics in this paper.
3.2 MVME Dataset Construction
Table 1: Departments distribution.